Lehrstuhl A für Mathematik Aachen, den 19.10.2001

Prof. Dr. R. L. Stens Dipl.-Math. B. Ohligs

1. Übung zu Zahlbereichserweiterungen

(Abgabe: Donnerstag, 25.10.2001, \underline{vor} der Übung)

Übungen: Die Übungsblätter sind auch im Internet erhältlich unter

http://www.mathA.rwth-aachen.de/lehre/WS01/Zbe

Die Homepage des Lehrstuhls A für Mathematik finden Sie unter

http://www.mathA.rwth-aachen.de

Scheinbedingungen: Maximal zwei Personen dürfen gemeinsam abgeben. Diplomkandidaten erhalten einen Übungsschein und Lehramtskandidaten einen qualifizierten Studiennachweis, falls mindestens 60% der Übungspunkte erreicht werden. Als Lehramtskandidat können Sie einen Leistungsnachweis erwerben, indem Sie \underline{zusätzlich} 1–2 Aufgaben in den Übungen vorrechnen.

Termine: Vorlesung: Mo 14.15 – 15.45 Uhr, Hörsaal IV, Mi 12.00 – 13.30 Uhr, Hörsaal IV Übung: Do 10.00 – 11.30 Uhr, Hörsaal V, \underline{Beginn 25.10.2001}

Aufgabe 1: Sei X eine nicht-leere Menge. Zeigen Sie, dass es keine surjektive Abbil-

dung von X auf die Potenzmenge P(X) von X gibt.  4

Aufgabe 2: Seien X, Y zwei nicht-leere Mengen, f : X → Y eine Abbildung und B ⊂ Y. Beweisen Sie:

f(f^{−1}(B)) = B ∩ f(X) ⊂ B.

Zeigen Sie anhand eines Gegenbeispiels, dass i.A. keine Gleichheit gilt. Finden Sie eine

hinreichende und notwendige Bedingung an die Abbildung f, so dass Gleichheit gilt.  4

Aufgabe 3: Seien X, Y zwei nicht-leere Mengen, f : X → Y eine Abbildung. Zeigen Sie die Äquivalenz der folgenden Aussagen:

(i) f ist injektiv,

(ii) f(A_{1} ∩ A_{2}) = f(A_{1}) ∩ f(A_{2}) (A_{1}, A_{2} ⊂ X),

(iii) f^{−1}(f(A)) = A (A ⊂ X),

(iv) f(A_{1} \ A_{2}) = f(A_{1}) \ f(A_{2}) (A_{1}, A_{2} ⊂ X).

6 Aufgabe 4: Seien X, Y zwei nicht-leere Mengen und f : X → Y eine Abbildung. Zeigen Sie:

a) f ist genau dann injektiv, wenn eine Abbildung g : Y → X mit g ∘ f = id_{X}

existiert. In diesem Fall ist jedes derartige g surjektiv.  2

b) f ist genau dann surjektiv, wenn eine Abbildung h : Y → X mit f ∘ h = id_{Y}

existiert. In diesem Fall ist jedes derartige h injektiv.  2

c) Sind Abbildungen g, h : Y → X mit g ∘ f = id_{X} und f ∘ h = id_{Y} gegeben, so

folgt g = h.  1

Man nennt g in a) ein Linksinverses und h in b) ein Rechtsinverses von f.

19

Georg Ferdinand Ludwig Philipp Cantor^{1}

Born: 3 March 1845 in St Petersburg, Russia Died: 6 Jan 1918 in Halle, Germany

Georg Cantor founded set theory and introduced the concept of infinite numbers with his discovery of cardinal numbers. He also advanced the study of trigonometric series.

Cantor attended the University of Zürich for a term in 1862 but then went to the University of Berlin where he attended lectures by Weierstrass, Kummer and Kronecker. He received his doctorate in 1867 from Berlin and accepted a position at the University of Halle in 1869, where he remained until he retired in 1913.

Cantor founded set theory and introduced the mathematically meaningful concept of infinite numbers with his discovery of transfinite numbers. He also advanced the study of trigonometric series and was the first to prove the nondenumerability of the real numbers.

His first papers (1870–1872) showed the influence of Weierstrass’s teaching, dealing with trigonometric series. In 1872 he defined irrational numbers in terms of convergent sequences of rational numbers. In 1873 he proved the rational numbers countable, i.e. they may be placed in 1–1 correspondence with the natural numbers.

Closely related to Cantor’s work in transfinite set theory was his definition of the continuum. Cantor’s work was attacked by many mathematicians, the attack being led by Cantor’s own teacher Kronecker. Cantor never doubted the absolute truth of his work despite the discovery of the paradoxes of set theory. He was supported by Dedekind, Weierstrass and Hilbert, Russell and Zermelo. Hilbert described Cantor’s work as the finest product of mathematical genius and one of the supreme achievements of purely intellectual human activity.

Cantor was given an honorary degree by the University of St Andrews in 1911. A major event planned in Halle to mark his 70th birthday in 1915 had to be cancelled because of the war. Cantor died in a psychiatric clinic in Halle in 1918.

1Aus: ‘The MacTutor History of Mathematics archive’ der University of St Andrews, Scotland.